A triangle has a permimeter of 161 miles. Each of the smaller two sides of the triangle is two-thirds the length of the longest side. Find the length of each side of the triangle.
Answer by checkley77(12844)

2X/3+2X/3+X=161
7X/3=161
7X=3*161
7X=483
X=483/769 IS THE LENGTH OF THE LOGEST SIDE.
2/3*69=46 FOR THE 2 SHORTER SIDES.
PROOF:
46+46+69=161
161=161